OpenStudy (anonymous):

Complete the square for the expression. Write the resulting expression as a binomial squared. x2 + 3x + _____

OpenStudy (anonymous):

x2 + 3x + 6 = (x + 3)2 x2 + 3x + = (x + )2 x2 + 3x + = (x + )2 x2 + 3x + 9= (x + 3)2

OpenStudy (unklerhaukus):

\[(x+a)^2=(x+a)(x+a)=x^2+xa+ax+a^2=x^2+2ax+a^2\]

OpenStudy (unklerhaukus):

if \(3x = 2ax\) what is \(a\)?